OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9 A.–9 + 2y B.–9 − 2y C.9 – 2y D.9 + 2y

OpenStudy (anonymous):

\[x=2y-9 --> -x+2y=9\] sub in this equation in 4x+5y=9 \[4(2y-9)+5y=9\] \[8y-36+5y=9\] \[13y=9+36\] \[13y=45\] \[y=\frac{ 45 }{ 13 }\] the answer is A

OpenStudy (anonymous):

(: